A father is 36 years older than his son. In 9 years, he will be 1/3 of his father's age. How old are they?
Answer by Maths68(1474)

Father’s age = F
Son’s age = S
Father is 36 years older. If we add 36 to the son’s age both will be the same age.
F=S+36 …………(1)
In 9 years means add 9 to both ages
F+9=S+9
Son will be 1/3 of his father’s age
(F+9)/3=S+9
F+9=3(S+9)
F+9=3S+27
F=3S+27-9
F=3S+18
Put the value of F from equation (1)
S+36=3S+18
36-18=3S-S
18=2S
18/2=S
S=9
Put the value of S in equation (1)
F=S+36
F=9+36
F=45
So Father is 45 years old and Son is 9 years old.